OpenStudy (anonymous):

If the domain is {0, 2, -6}, what is the range of y = -2x + 3? A) {0, 7, 29} B) {0, 7, 15} C) {3, -1, -9} D) {3, -1, 15}

OpenStudy (blurbendy):

substitute 0, 2, and -6 in for x in the equation y = -2x + 3 If you start with you 0 you get, y = -2(0) + 3 y = 0 + 3 y = 3 So 3 is in the range. Can you do the rest?

OpenStudy (anonymous):

Oh! Just substitute the domains for x. Easy! Thanks!! Answer is D. :)

OpenStudy (blurbendy):

Yup, good job.

OpenStudy (blurbendy):

Any other questions?

OpenStudy (anonymous):

No. But thanks for asking. :)

OpenStudy (blurbendy):

anytime